Bob is four more than three times the age of his little sister. If Bod is nineteen, then how old is little sister?
Answer by Cromlix(4381)

Hi there,
Bob is four more than three times
the age of his little sister.
B = 4 + 3S
If Bob is 19
4 + 3S = 19
3S = 19 - 4
3S = 15
S = 15/3
S = 5
Bob's sister is 5 years old.
hope this helps :-)